MSG Corporation issued $100,000 of 3-year, 6% bonds outstanding on December 31, 2020 for $106,000. The bonds pay interest annually and MSG uses straight-line amortization. On May 1, 2021, $10,000 of the bonds were retired at 112. As a result of the retirement, MSG will report:

Multiple Choice

  • a $600 loss.

  • a $667 loss.

  • a $1,200 loss.

  • a $1,200 gain.
Answer #1

Solution:

Value of $10000 bonds = 106000/100000*10000 = $10600

Premium = 10600 - 10000 = $600

Carrying value on May 1, 2021 = 10600 - 600*4/36 = $10533

Gain/ Loss on retirement = 10000*112%  - $10533 = $667 Loss

Option "$667 loss" is correct answer.